OpenStudy (anonymous):

Solve the equation for Y: 8x - 9y = 11 How do I solve this?

OpenStudy (anonymous):

get y and x alone

OpenStudy (welshfella):

first subtract 8x from both sides of the equation then divide both sides by -9

OpenStudy (anonymous):

8-8=0x=x 8-9=1y 8-11=3

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

Well, could I see the physical steps? telling me won't help unless I see how the steps are worked out.

OpenStudy (anonymous):

8x-9y=11 First you move the 8x because you have to leave the 9y alone 8x-9y=11 -8x -8x -9y=11-8x Now you have to move the -9, so the inverse of multiplying is dividing so you divide -9. -9y/-9= 11/-9 -8x/-9 since you can't have negative as a denominator it would be: y= -11/9+8x/9 You have to change the signs

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

make sense

OpenStudy (anonymous):

medal is best response